OpenStudy (supercait10):

the radius of a circular hat is 3.7 inches. Find the circumference use n=3.14 A 1.85 in B 11.618 in C 23.236 in D 221.2 in help plz medal and fan

OpenStudy (sleepyhead314):

Hello @supercait10 ! I saw your post on Twitter, so I came by to try to help :) The Equation for the Circumference of a Circle is 2 * n * (Radius) n = 3.14 (Radius) = 3.7 and we can just plug this into that equation :) Circumference of this Circle = 2 * (3.14) * (3.7) = [answer] I hope this helped! :) tag me if you need me to help clarify anything on this question ^_^

OpenStudy (supercait10):

so is it 22.2